OpenStudy (anonymous):

(x^3-8) is factored to (x-2)(x^2+2x=2) can someone explain this, or give me the formula to do a similar problem

OpenStudy (rsadhvika):

google a^3-b^3 formula

OpenStudy (anonymous):

why the b^3, the problem only has one exponent

OpenStudy (rsadhvika):

http://www.purplemath.com/modules/specfact2.htm

OpenStudy (rsadhvika):

observe that 8 = 2^3

OpenStudy (anonymous):

\[a^3-b^3 = (a-b)(a^2+ab+b^2)\]

OpenStudy (rsadhvika):

x^3-8 = x^3-2^3

OpenStudy (anonymous):

this x^3-8 = x^3-2^3 is the same as (x-2)(x^2+2x=2

OpenStudy (anonymous):

correct

OpenStudy (primeralph):

@Hero Let's try to arrive at the solution purely. Seems intricate.

OpenStudy (rsadhvika):

It is easy to see right, just 6 terms out of which 4 cancel out when you expand. you looking for derivation of formula ?

OpenStudy (anonymous):

@rsadhvika

OpenStudy (rsadhvika):

yeah ?

OpenStudy (primeralph):

Can someone post this as a new question? How to prove it.

OpenStudy (rsadhvika):

x^n-y^n = ... that proof you want ?

OpenStudy (primeralph):

Yes. For this case where n = 3.

OpenStudy (primeralph):

I'm working on it, but I'd like to see what others have.

OpenStudy (rsadhvika):

im doing the same right now, trying to understand the proof

OpenStudy (primeralph):

@Hero Got it.

OpenStudy (primeralph):